OpenStudy (anonymous):

how does 2x+y=1 and x+1=4-y

OpenStudy (anonymous):

Yo have to find x and y y = 1-2x y = 3-x y = y 1-2x = 3-x x = -2 y = 5

OpenStudy (anonymous):

They're two different equations.

OpenStudy (anonymous):

Yes, I know, but if you graph them, they intersect at a point and that point is where they have the same value

OpenStudy (anonymous):

The point connects the two equations together

OpenStudy (anonymous):

And that point is (-2,5) graphically